3,8,13,18,.........................,5008
What term is the number 5008?

Answer by stanbon(75887)

3,8,13,18,.........................,5008
What term is the number 5008?
----
The sequence is arithmetic with a = 3 and d = 8-3 = 5
----
Equation:
a + (n-1)d = 5008
3 + (n-1)5 = 5008
5n-5 = 5005
5n = 5010
n = 1002
===================
Cheers,
Stan H.
==============